# Entropy-conservative high-order methods for high-enthalpy gas flows

### v1

@@ 5 Conclusion
An approach to the computation of entropy-conserving fluxes for gases with arbitrary internal energies has been developed, based on the approximation of entropy via a numerical quadrature and subsequent application of the mean value theorem to express the jumps in entropy via jumps in the temperature. For a calorically perfect gas, the fluxes have been shown to reduce to well-known expressions available in literature. A significant advantage of the proposed approach is its independence on the exact expressions for the internal energy, as they are only required in the pre-computation step, and the flow simulations only perform linear interpolation, the cost of which is independent on the exact form of the internal energy.
The flux function has been implemented in the Trixi.jl framework and simulations of a Mach 10.5 flow over a cylinder were performed for three different internal energy functions. Comparisons with computations carried out using the DLR TAU solver show very good agreement in the flow-field quantities; with some discrepancies that can be attributed to a mismatch between the thermodynamical data used in the TAU solver and that used in the present work.
@@ Acknowledgments
This work has been supported by the German Research Foundation within the research unit DFG-FOR5409. Georgii Oblapenko also acknowledges the support of the Humboldt Foundation for his guest research stay at the German Aerospace Center (DLR) Institute of Aerodynamics and Flow Technology in Göttingen, during which the results with the DLR TAU solver had been obtained.

### v2

@@ 5 Conclusion
An approach to the computation of entropy-conserving fluxes for gases with arbitrary internal energies has been developed, based on the piece-wise linear approximation of the temperature-dependent internal energy and specific heat. For a calorically perfect gas, the fluxes have been shown to reduce to well-known expressions available in literature. A significant advantage of the proposed approach is its independence on the exact expressions for the internal energy, as they are only required in the pre-computation step, and the flow simulations only perform linear interpolation, the cost of which is independent of the exact expressions for the internal energy and specific heat capacities.
The flux function has been implemented in the Trixi.jl framework and simulations of a Mach 10.5 flow over a cylinder were performed for three different internal energy functions. Comparisons with computations carried out using the DLR TAU solver show excellent agreement in the flow-field quantities.
@@ Acknowledgments
This work has been supported by the German Research Foundation within the research unit DFG-FOR5409. Georgii Oblapenko also acknowledges the support of the Humboldt Foundation for his guest research stay at the German Aerospace Center (DLR) Institute of Aerodynamics and Flow Technology in Göttingen, during which the results with the DLR TAU solver had been obtained. The authors also thank Dr. Moritz Ertl of DLR Institute of Aerodynamics and Flow Technology Göttingen for his help with the DLR TAU simulations.